Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Lock acquisition: auxiliary lasers)
bersanetti, mantovani, ruggi, boldrini, pinto, tapia - 1:20 Saturday 20 February 2021 (50839) Print this report
ISC Shift: characterization of beatnote signals

The main points of today's shift are the following:

  • we carried on the characterization of the beatnote signals: we injected colored noise after the RFC error point (Sc_MC_RFC_Err_Pre, which is calibrated in µm), while the ALS was locked on both the arms, and the reallocation on the mirrors was active; the FlipMirror_ATR was up, to allow the presence of the beatnote signals. Noise was injected from 150 s starting from GPS = 1297763508; in Figure 1 there is the overall TF computed with this injection and the superposition of the swept sine from yesterday's shift;
  • during the shift we had an issue with the lock of the North Arm on the IR beam: a few seconds after the engagement of the lock, the corrections rapidly drifted away in a direction, causing the unlock; the problem disappeared after a while, but the reason is still unclear;
  • we did a test to apply an offset to a reallocation loop, which means setting an offset to the VCO correction, but only on the mirror part of the loop; we could see that only the main IR peak was scanned by the cavity (Figure 2); it is clear that in order to stay on the resonance, we need to have less noise on the reallocation loop or its equivalent for the coupled arms;
  • we set up most of the new code in LSC_Acl which will use the beatnote signals to lock CARM on the IMC:
    • we acquired the channels of the single-arm beatnote (NARM_BEAT_157MHz_DFREQ, WARM_BEAT_154MHz_DFREQ), the DC powers coming from CEB (NARM_BEAT_DC, WARM_BEAT_DC), the reflection DC signals from the External End Benches (ALS_NEB_PD_GREEN_DC, ALS_WEB_PD_GREEN_DC) and the reconstructed CARM/DARM beatnote signals (CARM_FREQ, DARM_FREQ); the LSC_ equivalents are not stored in the DAQ;
    • we took the chance of this edit to split the single, not stored ALS_NOISE channel used for the reallocation in two channels, one per arm, in order to resemble more the usual loop structure; the new channels will be stored in the DAQ;
    • we changed the way the CARM/DARM_LOCK_ON flags are generated, so that they can used when either both arms are locked on the IR (for backward compatibility) or both arms have the ALS loop reallocated on them; the new flags are stored in the DAQ as 1 Hz channels;
    • CARM_FREQ was added to the CARM_MC sensing matrix only, and DARM_FREQ to the DARM sensing matrix; the single-arm beatnote signals are added to the reallocation sensing matrix, which now can be filled with either the VCO error, the VCO correction (which is the usual sensing for this control) or the single-arm beatnote signal; the sensing matrix channels (to which is applied the setpoint) are called ALS_[N,W]Arm_RAW and they are saved in the DAQ;
    • given that we were at the end of the shift, we did not try to restart LSC_Acl, as all the changes need to be tested, and the last one will break the current code basis of the ALS Metatron nodes; this will be done next Monday afternoon.

PLEASE DO NOT RESTART OR DO A "RELOAD CONFIG" TO THE LSC_Acl PROCESS.

Images attached to this report
Comments to this report:
bersanetti, masserot - 1:00 Tuesday 23 February 2021 (50854) Print this report

Today, at the end of the morning shift, we addressed the pending changes to the LSC_Acl process:

  • we added the new channels we wanted to acquire in it to the Tolm packets coming from ALS_{NEB,WEB,CEB};
  • the LSC_Acl process could be then restarted; now we have available in LSC the reflected DC powers (ALS_*EB_GREEN_PD_DC_10KHz), the transmitted DC powers (ALS_CEB_*ARM_BEAT_DC), the single arms beatnote channels (ALS_CEB_*ARM_BEAT_15*MHz_dFreq) and the reconstructed CARM/DARM signals (ALS_CEB_*ARM_FREQ); since they are duplicate, the LSC_ equivalent are not stored in the DAQ; the relevant channels are already added to the sensing matrixes of interest;
  • now the reallocation sensing is not fixed anymore, but it is a ACL_SUM_CH weighted sum, where the *_Corr channel is the one currently in use; e.g.:
    • ACL_SUM_CH        ALS_NArm_RAW        "au"    1        0    ALS_NArm_INPUT    0    ALS_NArm_Corr    0    NARM_BEAT_157MHz_DFREQ
  • this change has been implemented in the Automation for the ALS system, but it hasn't been tested yet;
  • the logic for the lock flags for the CARM/DARM DOFs has been changed, in order to allow such lock only when, 1) both cavities are locked on the IR 2) both ALS loops are reallocated; the two conditions are treated with an "or" (not a custom "xor"), which should be enough given also the checks in Metatron; the new flags are on the DAQ.
Search Help
×

Warning

×